Onslow Co. purchased a used machine for $240,000 cash on January 2. On January 3, Onslow paid $8,000 to wire electricity to the machine and an additional $1,600 to secure it in place. The machine will be used for six years and have a $28,800 salvage value. Straight-line depreciation is used. On December 31, at the end of its fifth year in operations, it is disposed of.Journal entry worksheet 2 Record the first year year-end adjusting entry for the depreciation expense of the used machine. No
Answer #1

Answer:

Date Accounts titles and Explanation Debit ($) Credit ($)
Dec-31 Depreciation Expense           36,800
           Accumulated Depreciation -Machinery           36,800
(To Record the first year year-end adjusting entry for the depreciation expense)
Cost of machine = 240,000+8,000+1,600= $249,600
Depreciation (Straight-Line method) = Cost - Salvage / no of years
= (249,600-28,800)/6 yr =$36,800
Dec-31 Depreciation Expense           36,800
              Accumulated Depreciation -Machinery           36,800
( To Record the year of disposal year-end adjusting entry for the depreciation expense)